OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

Explain your answer so I could understand...

OpenStudy (phi):

Are you studying trigonometry? See http://www.khanacademy.org/math/trigonometry/v/basic-trigonometry for help on it. Also, it helps if you memorize sin(30º)= 1/2 cos(30º)= \(\sqrt{3}/2 \) and tan(30º)= sin(30º)/cos(30º)= \( 1/\sqrt{3} \)

OpenStudy (anonymous):

this is Pre-ALG

OpenStudy (phi):

People do study 30-60-90 triangles, because they have special properties. but part (c) in the question asks for sin,cos, tan, and that is trig. did you learn the definitions of sin,cos,tan?